OpenStudy (anonymous):

-7x-2y=-13 x-2y=11 Solve for both x and y I know how to do the problem so i just need to check my answer

OpenStudy (lgbasallote):

so what's your answer?

OpenStudy (anonymous):

Y= 4 and x= 19 what's your answer?

OpenStudy (lgbasallote):

hmm i dont think this is right

OpenStudy (lgbasallote):

if you substitute that to the first equation you'll have -7(19) -2(4) = -13 i doubt this results in an equality

OpenStudy (lgbasallote):

which method did you use? substitution or elimination?

OpenStudy (anonymous):

I used substitution and plug it in to the second equation

OpenStudy (lgbasallote):

okay let's try substitution from equation 2: x - 2y = 11 therefore x = 11 + 2y

OpenStudy (lgbasallote):

let's substitute this to the first equation -7(11 + 2y) -2y = -13 what do you get here?

OpenStudy (anonymous):

-77 -14y- 2y = 13

OpenStudy (lgbasallote):

right..now combine like terms

OpenStudy (anonymous):

16y= -13 +77 then add and divide so... Y= 4 like I said

OpenStudy (lgbasallote):

hmm what did you do...

OpenStudy (lgbasallote):

-77 - 14y - 2y = 13 -77 - 16y = 13 -77 - 13 = 16y

OpenStudy (lgbasallote):

try now

OpenStudy (lgbasallote):

lol..something's wrong here...wait a minute

OpenStudy (anonymous):

-30=6y soo.. Y=5?

OpenStudy (anonymous):

Ya the 13 is negative

OpenStudy (lgbasallote):

ohhhh lol -77 - 14y - 2y = -13 -77 - 16y = -13 -77 + 13 = 16y -64 = 16y -4 = y that's what you did wrong

OpenStudy (anonymous):

x=3 and y = -4

OpenStudy (anonymous):

How did u get -4

OpenStudy (lgbasallote):

-77 + 13 = -64

OpenStudy (lgbasallote):

-64/16 = -4

OpenStudy (anonymous):

x = 2y + 11 sub that into the first equation -7( 2y +11 ) -2y = - 13 -14y -77 - 2y = -13 -16y =64 y = -4

OpenStudy (lgbasallote):

do you get it now @azapata2 ?

OpenStudy (anonymous):

y = -4 x +8 = 11 x = 11 - 8 x = 3

OpenStudy (anonymous):

Ya thanks guys